OpenStudy (anonymous):

Write an equation that is parallel to 2x-3y=4 and crosses the point (-2,-4)?

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

parallel line means that they share the same slope

OpenStudy (anonymous):

i know there's a formula for this type of problem but i don't remember it..

OpenStudy (primeralph):

you've been given all you need.

OpenStudy (anonymous):

is it the point slope formula he posted?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use slope intercept to determine the slope of the line

OpenStudy (anonymous):

then use point slope in order to determine the equation of the parallel line that passes through the point

OpenStudy (anonymous):

okay! i can do that.:) thank you

OpenStudy (raden):

alternative : the equation of straight line which parallel to line ax+by=c and crosses the point (x1,y1) is ax + by = a(x1) + b(y1) so, it is 2x - 3y = 2(-2) - 3(-4) simplied, becomes 2x - 3y = 8